# I Learn Better If You Speak My Language: Understanding the Superior Performance of Fine-Tuning Large Language Models with LLM-Generated Responses

### v2

@@ 8 Conclusion
We systematically proved two hypotheses: LLMs are more familiar with responses generated by other LLMs when given similar inputs, and LLMs perform better on data they are familiar with. These hypotheses also explain why data annotated by LLMs yields better training outcomes than human-annotated data. The data generated by LLMs is more familiar to the trained LLMs, and training on unfamiliar data typically results in poorer outcomes. Based on this theory of familiarity, we have proposed several methods to generate data with high familiarity. Without using advanced LLMs, models can perform style transfers on ground truth to rewrite it in their own style. When using advanced LLMs for data processing, we can either use GPT-4 to read questions and generate answers in its own style, or employ our minimum change data generation method, which in some tasks, yields even better training results than direct responses from GPT-4.
@@ 9 Limitations
Perplexity can serve as a straightforward tool for measuring familiarity, but it is not very precise. It only provides relatively accurate predictions of performance decline when there are substantial differences in perplexity levels. We need a better tool to predict a model’s familiarity with a dataset and how that familiarity impacts the model’s learning.
There may be additional forms of familiarity, such as linguistic familiarity and logical familiarity. Accurately identifying and measuring the impact of these forms of familiarity presents a challenge.
Appendix A Appendix: Comparing varients of GPT-4 created training datasets
@@
GPT-4 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4 for answers. For classification tasks, we provide GPT-4 with the gold-label (which does not include human-annotated rationales) and the input, and then use its responses. We combine these answers with the original questions to create the GPT-4 Answer Training Dataset. We purposefully do not supply GPT-4 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes. If these processes were included, it could lead to a portion of GPT-4’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
Step by step transformation on GT: For classification tasks, we provide GPT-4 with both the questions and the ground truth target labels along with human-annotated rationales. For mathematical problems, we present the problems along with their answers, including the solutions process. We instruct GPT-4 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps. If the existing content is already detailed enough, no additional details are added. We also instruct GPT-4 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
Detailed Step-by-Step Transformation of Ground Truth: This data construction method is very similar to the Step by Step method. The only difference is that we specifically instruct GPT-4 to add extra details when the original ground truth lacks sufficient information.
Add details to GT: This approach involves enhancing the existing ground truth data by supplementing it with additional details to provide a richer context when necessary. We direct GPT-4 to enhance the content while maintaining the inherent language style of the original data. This method is applied selectively, only when the original information lacks sufficient clarity or depth to support effective model training and understanding.
@@
Rewrite GT in GPT-4 style words: For this method, we present GPT-4 with the ground truth data, including human-annotated rationales and problem-solving steps. The objective is to have GPT-4 rewrite the ground truth using its own language style.
Datasets with explicit reasoning
Math Commensense Reasoning NLI
Method Training Dataset and Model Type GSM8K Math Algebra ECQA Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 4.72 128.882
GPT-4 Answer Directly 0.597 0.246 0.597 2.34 179.941
Rewrite GT in GPT-4 style words 0.471 0.192 0.6 4.77 156.696
Step by step 0.574 0.224 0.591 2.27 282.9
Add details to GT 0.59 0.235 0.596 3.80 272.844
Step by step transformation of GT 0.499 0.190 0.604 2.67 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.506 0.196 0.602 2.59 315.028
Groundtruth GSM8K, Llama 0.364 0.141 0.565 3.27 128.882
GPT-4 Answer Directly 0.428 0.127 0.575 2.30 179.941
Rewrite GT in GPT-4 style words 0.394 0.127 0.58 3.65 156.696
Step by step 0.402 0.134 0.553 2.19 290.2
Add details to GT 0.38 0.128 0.579
Step by step transformation of GT 0.394 0.146 0.573 2.42 235.892
Detailed Step-by-Step Transformation of Ground Truth 0.396 0.114 0.553 2.37 315.028
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 5.93 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 2.20 328.764
Step by step 0.533 0.239 0.583 2.25 452
Add details to GT 0.445 0.230 0.57 3.148 417.35
Step by step transformation of GT 0.444 0.210 0.577 2.65 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.468 0.223 0.582 2.62 484.368
Groundtruth Math algebra, Llama 0.36 0.126 0.575 5.17 163.355
GPT-4 Answer Directly 0.35 0.150 0.561 2.591 328.764
Step by step 0.344 0.144 0.593 2.55 450.1
Add details to GT 0.38 0.128 0.579 3.31 417.35
Step by step transformation of GT 0.376 0.140 0.565 2.80 370.349
Detailed Step-by-Step Transformation of Ground Truth 0.299 0.110 0.545 2.816 484.368
Groundtruth ECQA, Mistral 0.258 0.134 0.68 51.28 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 6.35 176.536
Rewrite GT in GPT-4 style words 0.384 0.194 0.721 9.26 134.338
Step by step 0.481 0.203 0.71 4.66 325.2
Add details to GT 0.422 0.204 0.731 9.08 238.292
Step by step transformation of GT 0.462 0.194 0.681 6.858 218.74
Detailed Step-by-Step Transformation of Ground Truth 0.487 0.186 0.68 5.61 322.855
Groundtruth ECQA, Llama 0.132 0.079 0.631 12.16 73.508
GPT-4 Answer Directly 0.379 0.155 0.656 5.33 176.536
Rewrite GT in GPT-4 style words 0.282 0.115 0.664 6.84 134.338
Step by step 0.363 0.140 0.648 4.46 337.3
Add details to GT 0.062 0.097 0.663 8.47 238.292
Step by step transformation of GT 0.363 0.139 0.648 5.54 218.748
Detailed Step-by-Step Transformation of Ground Truth 0.135 0.106 0.66 5.08 322.855
Table 6: Appendix A: Additional experiments for Table2
Appendix B Appendix: Minimum Change Vs GPT4 answer
Datasets with explicit reasoning
Math Commensense Reasoning NLI Code
Method Training Dataset and Model Type GSM8K Math Algebra ECQA ESNLI HumanEval MBPP Perplexity Toekn Count
Groundtruth GSM8K, Mistral 0.434 0.162 0.594 0.597 4.53 128.88
GPT-4 Answer Directly 0.597 0.246 0.597 0.587 2.34 179.941
Mistral Minimum Change Data 0.562 0.234 0.597 0.558 1.67 177.156
llama minimum change data 0.547 0.242 0.621 3.19 166.152
Groundtruth GSM8K, Llama 0.371 0.143 0.566 0.559 3.27 130.856
GPT-4 Answer Directly 0.428 0.127 0.575 0.568 2.30 183.852
Llama Minimum Change Data 0.433 0.140 0.602 0.602 2.22 166.15
Mistral Minimum Change Data 0.402 0.141 0.576 2.249 177.156
Groundtruth Math algebra, Mistral 0.264 0.206 0.554 0.574 5.83 163.355
GPT-4 Answer Directly 0.553 0.301 0.608 0.555 2.19 328.764
Mistral Minimum Change Data 0.536 0.313 0.622 0.558 1.85 297.609
llama minimum change data 0.546 0.296 0.63 2.47 269.688
Groundtruth Math algebra, Llama 0.36 0.126 0.575 0.552 5.14 162.417
GPT-4 Answer Directly 0.35 0.150 0.561 0.564 2.59 326.573
Llama Minimum Change Data 0.412 0.166 0.602 0.554 2.268 269.68
Mistral Minimum Change Data 0.32 0.160 0.568 2.53 297.609
Groundtruth ECQA, Mistral 0.258 0.1348 0.68 0.544 51.29 73.508
GPT-4 Answer Directly 0.462 0.223 0.722 0.589 6.35 176.53
Mistral Minimum Change Data 0.433 0.192 0.699 0.542 3.798 112.591
LLama Minimum Change Data 0.483 0.212 0.709 4.89 128.695
Groundtruth ECQA, Llama 0.132 0.079 0.631 0.431 11.95 76.118
GPT-4 Answer Directly 0.379 0.155 0.656 0.479 5.33 183.182
Llama Minimum Change Data 0.392 0.128 0.649 0.522 3.619 128.695
Mistral Minimum Change Data 0.28 0.126 0.647 5.58 112.591
Groundtruth ESNLI, Mistral 0.417 0.172 0.594 0.814 378.7 29.925
GPT-4 Answer Directly 0.464 0.214 0.592 0.804 6.54 137.967
Mistral Minimum Change Data 0.453 0.206 0.602 0.831 3.38 82.367
Llama Minimum Change Data 0.432 0.203 0.63 0.836 4.15 97.078
Groundtruth ESNLI, Llama 0.332 0.125 0.582 0.751 1201.70 29.963
GPT-4 Answer Directly 0.386 0.123 0.564 0.776 5.28 143.739
Llama Minimum Change Data 0.357 0.128 0.578 0.796 3.74 97.078
Mistral Minimum Change Data 0.346 0.130 0.546 0.811 4.34 82.367
Groundtruth HumanEval, Mistral 0.362 0.191 0.583 0.531 0.323 5.69 71.60
GPT-4 Answer Directly 0.313 0.162 0.581 0.442 0.365 4.71 107.9
Mistral Minimum Change Data 0.4055 0.202 0.5915 0.541 0.408 2.21 109.073
Llama Minimum Change Data 0.4115 0.179 0.5835 0.5505 0.426 2.451 122.0
Groundtruth HumanEval, Llama 0.0705 0.083 0.528 0.339 0.146 5.75 70.8
GPT-4 Answer Directly 0.1245 0.105 0.553 0.473 0.158 4.33 107.14
Llama Minimum Change Data 0.2445 0.389 0.5705 0.5175 0.213 2.21 122.0
Mistral Minimum Change Data 0.131 0.103 0.528 0.4235 0.183 2.75 109.07
Groundtruth MBPP, Mistral 0.392 0.175 0.527 0.481 0.276 7.78 71.60
GPT-4 Answer Directly 0.399 0.186 0.568 0.553 0.354 3.419 786
Mistral Minimum Change Data 0.405 0.179 0.601 0.48 0.354 2.817 63.86
Llama Minimum Change Data 0.332 0.179 0.56 0.511 0.364 4.55 54.96
Groundtruth MBPP, Llama 0.328 0.131 0.556 0.537 0.2 4.74 73.90
GPT-4 Answer Directly 0.351 0.139 0.574 0.553 0.202 3.33 69.78
Llama Minimum Change Data 0.34 0.130 0.569 0.539 0.204 3.44 54.96
Mistral Minimum Change Data 0.333 0.123 0.563 0.544 0.218 3.45 63.86
Table 7: Appendix B: More results for Table 5. Comparing the experimental results of GPT4, groundtruth and minimum change
Appendix C Appendix: Minimum Change Prompt and Method
Figure 3: Appendix C: Minimum Change Data Correction Examples
@@ Appendix C Appendix: Minimum Change Prompt and Method
  • Minimal Modification uses GPT-4 to adjusts the initial prediction with the fewest possible changes to correct errors, deliberately preserving the original distribution by keeping as many words unchanged as possible. An illustration of such a modification is provided in Figure 2. We have detailed the prompt used to guide GPT-4 in the ’Minimum Change Prompt’ section of the Appendix.
@@
  • Training with Modified Predictions the minimum changed predictions are subsequently used as target labels for model training.
Figure 4: Appendix C: Minimum Change Prompt Example
@@
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ’Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
Appendix D Appendix: LLaMA groundtruth style transfer failure example

### v3

@@ 7 Conclusion
We conducted a series of experiments involving the rewriting of target responses to validate our two proposed hypotheses: 1) LLMs are more familiar with responses generated by themselves or other LLMs when given the same question; 2) LLMs exhibit improved performance when trained on data with which they are already familiar. These hypotheses help explain why data annotated by LLMs leads to superior training outcomes compared to human-annotated data on reasoning tasks, from a familiarity perspective. These findings suggest new directions for optimizing data annotation processes to enhance model performance.
@@ 8 Limitations
The effectiveness of the ground truth transformation method and the minimum change method can be limited by their accuracy. When an LLM is not powerful enough, it may not be able to successfully rewrite the ground truth in its own style while maintaining the correctness of the ground truth. Additionally, when performing minimal changes to a model’s initial prediction, we notice that GPT-4 often fails to correct the initial prediction accurately. Although GPT-4 may correct some errors, the minimally changed response can still be incorrect. Furthermore, minimal change is a challenging task. Achieving such a process requires an advanced LLM, as the LLM needs to fully understand the solution and follows the “minimal change” principal.
Appendix A Comparing varients of GPT-4/Claude 3.5 created training datasets
We instruct GPT-4/Claude 3.5 to generate target responses in various ways, and use these labels for training. The training results are summarized in Table 6 and Table 7, and we describe the different methods used to instruct GPT-4/Claude 3.5 to generate target responses.
@@
GPT-4/Claude 3.5 Answer Directly: For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4/Claude 3.5 for answers. For classification tasks, we provide GPT-4/Claude 3.5 with the gold-label (which does not include human-annotated rationales) and the input questions, and then use its responses. We combine these answers with the original questions to create the GPT-4/Claude 3.5 Answer Training Dataset. We purposefully do not supply GPT-4/Claude 3.5 with the solutions or rationales for the math, coding, or classification problems to prevent it from merely replicating the ground truth’s problem-solving and analytical processes. If these processes were included, it could lead to a portion of GPT-4’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Response (No GT) response.
"""We have the {question}
1. We wish you to answer the question.
2. You must answer the question (with inference process) directly without say anything else. Please not saying anything 'like sure I can help you with' or 'sure, i will not mention the gold label'
3. You will inference first then put the Final Answer (NUMBER_HERE) at the end of the prediction like this
INFERENCE HERE
Final Answer: NUMBER_HERE"""
@@
Rewrite Ground Truth: For this method, we present GPT-4/Claude 3.5 with the ground truth data, including human-annotated rationales and problem-solving steps. The objective is to have GPT-4/Claude 3.5 rewrite the ground truth using its own language style.
The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Response (Rewrite GT) response.
"""Given the question: {question}
and the groundtruth: {groundtruth}
Please states the prediction in your own words. The groundtruth is 100% correct. You should not change the problem solving logic of the groundtruth. just restates it in your own words.
1. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
2. (important format) You must generate the groundtruth directly. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
3. (important format) Please make sure the Final Answer: {gold_label} is placed at the end of the modified prediction."""
Step-by-step: We specificly tell GPT-4/Claude 3.5 to solve the problem step by step. For mathematics and coding problems, we directly present the problems from the training dataset to GPT-4/Claude 3.5 for answers. For classification tasks, we provide GPT-4/Claude 3.5 with the gold-label (which does not include human-annotated rationales) and the input questions, and then use its step by step responses. We combine these answers with the original questions to create the GPT-4/Claude 3.5 Step-by-Step Response (No GT) Dataset. We purposefully do not supply GPT-4/Claude 3.5 with the solutions or rationales for the math, coding, or classification problems to prevent it from replicating the ground truth’s problem-solving and analytical processes. If these processes were included, it could lead to a portion of GPT-4/Claude 3.5’s predictions not being generated in its own style, thus rendering our perplexity measurements—which assess how well an LLM handles predictions from other LLMs—less accurate.
The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Response (No GT) response.
We have the question and the groundtruth. Please reformat the groundtruth in step by step manner with details.
Question: {question}
Groundtruth: {groundtruth}
1. We wish you to regenerate a new groundtruth. The new groundtruth solve the problem step by step. If you believe the groundtruth is not detail enough, you could add details.
2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
3. (important format) You must generate the groundtruth with the step by step inference process directly. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
4. (important format) You will inference first then put the Final Answer: {gold_label}
at the end like this
INFERENCE HERE
Final Answer: {gold_label}
Step-by-Step Transformation of GT: For classification tasks, we provide GPT-4/Claude 3.5 with both the questions and the ground truth target labels along with human-annotated rationales. For mathematical problems, we present the problems along with their answers, including the solutions process. We instruct GPT-4/Claude 3.5 to rewrite the human-annotated solutions or rationales in a step-by-step format, adding details where necessary to fill any gaps. If the existing content is already detailed enough, no additional details are added. We also instruct GPT-4/Claude 3.5 to preserve the original linguistic style as closely as possible by frequently incorporating language from the original ground truth.
The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Response (Rewrite GT) response.
"""We have the question and the groundtruth. Please reformat the groundtruth in step by step manner. You must try your best to keep the original words and logic unchange. You are reformating the groundtruth into step by step solution. you could add new words, but you are trying to keep original words from groundtruth unchanged.
Question: {question}
Groundtruth: {groundtruth}
1. We wish you to reformat a new groundtruth. The new groundtruth are reformated a new groundtruth which solve the problem step by step. you could add new words, but you are trying to not change the original words.
2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
3. When adding step by step inference process, you should imitate the language styles of the original groundtruth.
4. (important format) You must generate the groundtruth with the step by step inference process directly. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
5. (important format) You will inference first then put the Final Answer: {gold_label}
at the end like this
INFERENCE HERE
Final Answer: {gold_label}"""
Detailed Step-by-Step Transformation of GT: This data construction method is very similar to the Step by Step method. The only difference is that we specifically instruct GPT-4/Claude 3.5 to add details when the original ground truth lacks sufficient information.
The following prompt instructs GPT-4/Claude 3.5 to create the GPT-4/Claude 3.5 Step-by-Step Detailed Response (Rewrite GT) response.
"""We have the question and the groundtruth. Please reformat the groundtruth in step by step manner with details for each step. When adding detail, you must try your best to keep the original words and logic unchange. You are reformating the groundtruth with more details on each step instead of rewrite the groundtruth, but you are trying to keep original words from groundtruth unchanged.
Question: {question}
Groundtruth: {groundtruth}
1. We wish you to reformat a new groundtruth. The new groundtruth are reformated a new groundtruth which solve the problem step by step with detail on each step.
2. You will pretend as you do not know the groundtruth, because we will use your prediction as target labels to train our model.
3. When adding step by step inference process, you should imitate the languaguage styles of the original groundtruth.
4. (important format) You must generate the groundtruth with the detailed step by step inference process directly. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
5. (important format) You will inference first then put the Final Answer: {gold_label}
at the end like this
INFERENCE HERE
Final Answer: {gold_label}"""
GPT-4/Claude 3.5 Answer with Higher Perplexity: In this method, we first create a ‘paraphrased GPT-4/Claude 3.5 answer’ dataset. Specifically, we provide GPT-4 with the questions and the target responses from the ‘GPT-4/Claude 3.5 Response (No GT)’, instructing it to paraphrase only the target responses. We include the questions in the paraphrasing instruction prompt to ensure that the keywords in the paraphrased answers match those in the questions. This approach helps maintain content relevance. We then measure the perplexity of the ‘paraphrased GPT-4/Claude 3.5 response’ on the model that will be trained. We compare the perplexity of the target responses from the ‘GPT-4/Claude 3.5 Response (No GT)’ to the target responses from the ‘paraphrased GPT-4 response’. We select the one with the higher perplexity and collect them paired with the original question to form a new dataset, named ‘GPT-4/Claude 3.5 Answer with Higher Perplexity’.
The following prompt instructs GPT-4/Claude 3.5 to create the paraphrased response.
"""Given the question: {question}
and the prediction: {gpt4_prediction}
1. Please paraphrase the prediction. The prediction might be wrong sometimes, but you do not need to correct it. just paraphrase it.
2. (important format) You must create the paraphrased prediction directly without say anything else. Please not saying anything like 'sure I can help you with' or 'sure, i will not mention the gold label'
4. (important format) Please make sure the Final Answer: {gold_label_type} is placed at the end, which means you will not paraphrase the phrase 'Final Answer'."""
Training Data Construction for Coding Tasks
When constructing training datasets for coding tasks (such as HumanEval and MBPP), we follow a specialized pipeline. To build training datasets generated by models like GPT-4,Claude 3.5, Llama2, or Mistral, we first feed the problem to an LLM to generate an initial prediction. Next, we use GPT-4 to extract the corresponding code. These extracted codes are then assembled into target responses. The specific code extraction prompt is detailed below.
f"""You need to extract the code from the previous prediction. The prediction answer the QUESTION.
QUESTION: {modified_question}
Previous prediction: {previous_prediction}
Explaination of extract code:
1. You have to extract code from the previous prediction. Your answer will be evaluated to see if it will pass the test, thus of course you need to extract the code and comments without including any other words.
2. You only need to extract the final code for the fuction for the test case{item['test_list'][0]}. You do not care about the test cases.
3. Please directly show me the extracted code with no other words. We will run the code to test if it is correct, thus do not extract other words.
4. you suppose to extract the code from the previous prediction. You should never correct the code on your own even if it is wrong.
Appendix B Minimum Change Vs GPT4 answer
Table 8 summarizes the performance differences between GPT-4 and Minimum Change. According to the table, models trained on Minimum Change predictions perform comparably to those trained on GPT-4 generated predictions. The average length of the Minimum Change predictions is noticeably shorter than that of GPT-4 generated predictions for most tasks. Please note that Minimum Change is designed to maximize familiarity for the target model. For example, when testing on the HumanEval dataset, the cross-domain and in-domain performance of Llama’s Minimum Change data is noticeably higher than Mistral’s Minimum Change data when the target model is Llama. This is because Llama’s Minimum Change data focuses primarily on maximizing familiarity specifically for the Llama model.
Appendix C Minimum Change Prompt and Method
The minimum change pipeline is illustrated in Figure 3. The prompt used to instruct GPT-4 for creating the minimum change response is illustrated in Figure 4.
@@ Appendix C Minimum Change Prompt and Method
  • Minimal Modification uses GPT-4 to adjusts the initial prediction with the fewest possible changes to correct errors, deliberately preserving the original distribution by keeping as many words unchanged as possible. An illustration of such a modification is provided in Figure 3. We have detailed the prompt used to guide GPT-4 in the ‘Minimum Change Prompt’ section of the Appendix.
@@
  • Training with Modified Predictions the minimum changed predictions are subsequently used as target responses for model training.
@@
This method involves minimal intervention by GPT-4 in the initial predictions to ensure that the original language style is largely retained. To effectively guide GPT-4 in producing these minimally changed training datasets, we used a specific prompt, outlined in the ‘Minimum Change Prompt’ section of the Appendix, which includes two to three examples of minimal changes along with explanations for each modification.
Appendix D Llama2 groundtruth style transfer failure example